# Searching for quantum-gravity footprint around stellar-mass black holes

### v1

@@ Abstract
According to the asymptotically safe gravity, black holes can have characteristics different from those described according to general relativity. Particularly, they are more compact, with a smaller event horizon, which in turn affects the other quantities dependent on it, like the photon ring and the size of the innermost stable circular orbit. We decided to test the latter by searching in the literature for observational measurements of the emission from accretion disk around stellar-mass black holes. All published values of the radius of the inner accretion disk were made homogeneous by taking into account the most recent and more reliable values of mass, spin, viewing angle, and distance from the Earth. We do not find any significant deviation from the expectations of general relativity. Some doubtful cases can be easily understood as due to specific states of the object during the observation or instrumental biases.
@@ 1 Introduction
The description of all physical phenomena in the Universe in one unified theory is the ultimate dream of theoretical physics. Despite of decades of theoretical research [1, 2], little has been obtained on the observational/experimental side. Any experimental or observational evidence of quantum gravity remains elusive still today [3]. Today more than ever, it is necessary to find some observational constrain, to try stopping the uncontrolled proliferation of theories.
Among the different available variants of quantum gravity theories, asymptotically safe gravity (ASG) could offer some interesting opportunities. First developed by Steven Weinberg [4], it has its pillars on quantum field theory and renormalization group [5, 6]. Most important, it is expected to have observational consequences on astrophysical and cosmological scales [7], and in particular around black holes [8, 9, 10, 11]. For large values of the distance $r>>l_{\rm Planck}$, where $l_{\rm Planck}\sim 1.62\times 10^{-33}$ cm is the Planck length, the gravitational constant $G(r)$ becomes [9]:
$G(r)\sim G_{\rm N}\left(1-\frac{\xi}{r^{2}}\right)$ (1.1)
where $G_{\rm N}\sim 6.67\times 10^{-8}$ cm3 g-1 s-2 is the classical gravitational constant, and $10<\xi<$ 1 is the coupling coefficient.
The main effect on black holes is a more compact object, although without curvature singularities [10]. A smaller event horizon affects also all the other related quantities such as the photon ring [10], the ergosphere [9] and the radius of the innermost stable circular orbit $r_{\rm isco}$ [11]. The effects of ASG should be larger for smaller masses, although the constraint of $10<\xi<$ 1 sets severe limits on them. When the mass of the compact object is much larger than the Planck mass – which is any real case – the critical surfaces of an ASG black hole should not be distinguishable from the classical one according to [8] (see their Fig. 4 and 5). However, [11] suggested that differences up $\sim 30$% in the $r_{\rm isco}$ are expected in the case of maximally rotating prograde black holes already with a small value of $\xi$ (see his Table I). Moreover, the constraint on $\xi$ was put to question in Eichhorn and Held [10], who claim that values greater than 1, which lead to larger deviations from GR, can be admitted within the framework of the ASG. The same authors also stress that “irrespective of theoretical considerations, any observational avenue to put constraints on deviations from GR, should be explored” [10]. This is also our philosophical approach to the problem, searching for any deviation from GR around black holes.
@@
Some attempts have been already made based on X-ray spectroscopy [12, 13, 14], but to date no deviations were found. We focus our study on the search for deviations from the value of the $r_{\rm isco}$ in the case of stellar-mass black holes calculated according to general relativity. Since these objects are the smallest black holes (primordial black holes were never observed to date), the effects of ASG – if any – should be maximized. Anyway, it should be noted that the theoretical minimum mass of a black hole is estimated to be $\sim 2.33M_{\odot}$ [15], which is still $\sim 38$ orders of magnitude greater than the Planck mass (but remind the above quote by Eichhorn and Held [10]). In addition, since the peak of the emission from the accretion disk is inversely proportional to the square root of the mass of the compact object, it falls in the X-ray energy band, where are less problems of contamination from the nearby environment. In the case of supermassive black holes, the disk emission peaks in the ultraviolet, where there are contaminations from emission lines of the broad-line region and the host galaxy.
As a first step, since we need of observations of black holes in specific states (high/soft, see Sect. 3), we searched in all the available literature for measurements of the inner radius of the accretion disk. Then, we made all the values homogeneous to compare each other.
@@ 2 A short summary of the accretion disk theory and models
According to the standard theory [16, 17], the accretion disk around black holes is optically thick and geometrically thin, although there are other models, particularly in the case of very high or very low mass accretion rate [18]. From the observational point of view, the disk can be considered as a series of rings of increasing size going outward, each one emitting black body radiation with increasing temperature $T(r)$ as the radial distance $r$ from the black hole decreases (e.g. see Fig. 2 in [19]) according to:
$T(r)\propto r^{-\frac{3}{4}}$ (2.1)
This model is named multicolor accretion disk [20, 21, 22], and is implemented as diskbb model in the xspec software package [^2] https://heasarc.gsfc.nasa.gov/xanadu/xspec/. The highest temperature $T_{\rm in}$ corresponds to the innermost part of the disk, the one closest to the event horizon.
The disk luminosity $L_{\rm d}$ [erg s-1] can be calculated via the Stefan-Boltzmann law:
$L_{\rm d}=4\pi r_{\rm in}^{2}\sigma T_{\rm in}^{4}$ (2.2)
where $\sigma=5.67\times 10^{-5}$ erg cm-2 s-1 K-4 is the Stefan-Boltzmann constant, and $T_{\rm in}$ [K] is the innermost temperature at the radius $r_{\rm in}$ [cm]. This equation is valid when the outer radius of the accretion disk $r_{\rm out}$ is much larger than $r_{\rm in}$, so that the outer temperature $T_{\rm out}$ can be considered negligible.
It is worth noting that X-ray spectra cannot measure the innermost effective temperature, because of Comptonization. The observed color temperature $T_{\rm col}$ is greater than the effective one by a factor $\kappa\sim 1.7-2.0$, which depends on the accretion rate [23]. This affects also the effective location of the inner radius, which should be corrected by a factor $\varsigma=0.412$ [24]. Therefore, to use the observed quantities, Eq. (2.2) is rearranged as:
$L_{\rm d}=4\pi\sigma\left(\frac{r_{\rm in}}{\varsigma}\right)^{2}\left(\frac{T_{\rm col}}{\kappa}\right)^{4}$ (2.3)
@@
$L_{\rm d}=2\pi\frac{d^{2}}{\cos i}F_{\rm bol}$ (2.4)
where $F_{\rm bol}$ is the disk bolometric flux [erg cm-2 s-1], $d$ is the luminosity distance [cm], and $i$ is the viewing angle of the disk [deg]. By combining Eqs. (2.3) and (2.4), it is then possible to estimate the inner radius of the accretion disk:
$r_{\rm in}=\varsigma\kappa^{2}d\sqrt{\frac{F_{\rm bol}}{2\cos i\sigma T_{\rm col}^{4}}}$ (2.5)
The normalization $N$ of the diskbb model – that can be estimated via spectral fitting – includes some of the above factors, and is equal to:
$N=\left(\frac{r_{\rm col}}{d_{10}}\right)^{2}\cos i$ (2.6)
where $d_{10}$ is the luminosity distance in units of $10$ kpc, and $r_{\rm col}=r_{\rm in}/\varsigma$ is the color radius, which must be corrected according to the above cited factors as:
$r_{\rm in}=\varsigma\kappa^{2}r_{\rm col}=\varsigma\kappa^{2}d_{10}\sqrt{\frac{N}{\cos i}}$ (2.7)
Despite its simplicity, the diskbb model performs reasonably well, as it results from the comparison with more detailed models [25, 19, 26, 27], although with some caveats [28, 19].
@@ 3 Gathering data
Eqs. (2.5) and (2.7) offer two possibilities to estimate the innermost radius of the accretion disk. In addition to the quantities measured via the fit of the X-ray spectrum ($N$, $T_{\rm col}$, $F_{\rm bol}$), it is necessary to know the distance $d$, the viewing angle $i$, but also the mass of the black hole $M$ and its spin $a$ to estimate the radius of the innermost stable circular orbit $r_{\rm isco}$ according to GR (Kerr metric, [17]):
$r_{\rm isco}=r_{\rm g}\left[3+Z_{2}-\sqrt{(3-Z_{2})(3+Z_{1}+2Z_{2})}\right]$ (3.1)
@@
The values of $r_{\rm g}$, $a$, $d$, and $i$ changed often during time, depending on the instrument evolution and the observational constrains (e.g. finding the stellar companion is often difficult, given the presence of many molecular clouds and interstellar dust). We adopted the latest values, which are not always consistent with those adopted by many other authors when converting their observational data. For example, many authors did not report the normalization $N$ of the diskbb model, but only the calculated $r_{\rm in}$ or even $r_{\rm col}$. Therefore, to extract a value comparable with other observations made at different epochs and with different instruments, it is necessary to know which values of $d$ and $i$ were adopted in the conversion. This is not granted: sometimes the values used in the conversion were not written in the paper, which made it impossible to recalculate the original normalization. Other problems encountered when searching in the literature were: missing measurement errors, missing units of measurement, plain errors and/or typos. In case of errors, it was often possible to correct them by recalculating the described model. For example, when finding a weird normalization $N$, it is possible to recalculate it if $T_{\rm col}$ and $F_{\rm bol}$ were reported. Fluxes were often reported without measurement errors: in this case, a typical $10$% error was adopted.
Yet another issue regards the instruments. In the case of stellar-mass black holes, the color temperature of the multicolor accretion disk peaks in the range $\sim 0.1-1$ keV. However, many detectors have a low-energy threshold of a few keV, which made it difficult to correctly extrapolate to lower energies. For example, one of the most used and prolific satellites in the observation of X-ray binaries was the Rossi X-ray Timing Explorer (RXTE) [^3] https://heasarc.gsfc.nasa.gov/docs/xte/xtegof.html. The Proportional Counter Array (PCA) worked in the $2-60$ keV energy range, and its low-energy threshold degraded up to $3$ keV with the age. The $2$ keV threshold makes it difficult to correctly extrapolate a $T_{\rm col}=0.4$ keV, because only a small part of the spectrum can be detected, as shown in Fig. 1.
@@
In addition to the above cited problems, it is necessary to take into account that the accretion disk does not extend always down to the innermost stable circular orbit. Stellar-mass black holes change their activity from high/soft to low/hard states, and vice versa, passing through some intermediate states, depending on the accretion rate [37, 38, 39, 40]. High/soft states occur at high accretion rates. The X-ray spectrum is dominated by the thermal emission of a geometrically-thin optically-thick accretion disk, which can extend down to $r_{\rm isco}$, and it accounts for more than $50$% of the bolometric flux [38]. At low accretion rates (low/hard states), the thin disk is weak and cool, and a mildly relativistic radio jet is present. The X-ray spectrum is dominated by the optically-thin emission from the corona, modelled with a hard power law or a more detailed Comptonization model. The intermediate state is an obvious mixture of the two above cited states. There are also many other variations, distinguished by small details [^4] See the plot: https://www.issibern.ch/teams/proaccretion/Documents.html. However, for our purposes, we focus our attention on high/soft states, when the accretion disk extends downward the innermost stable circular orbit. This is also a requirement to have a reliable measurement of $r_{\rm in}$, because the multicolor disk model tends to underestimate the inner radius when the corona significantly contributes to the bolometric flux [28], as it happens in the intermediate states, for example.

### v2

@@ Abstract
According to the asymptotically safe gravity, black holes may have characteristics different from those described according to general relativity if the running of the gravitational constant coupling happens at low energies. Particularly, they should be more compact, with a smaller event horizon, which in turn affects the other quantities dependent on it, like the photon ring and the size of the innermost stable circular orbit. We decided to test the latter hypothesis by searching in the literature for observational measurements of the inner radius of the accretion disk around stellar-mass black holes. We selected the smallest values measured when the disk was in high/soft state, made them homogeneous by taking into account the most recent and more reliable values of mass, spin, viewing angle, and distance from the Earth, and compared with the expectations of the Kerr metric. We do not find any significant deviation. Some doubtful cases can be easily understood as due to specific states of the object during the observation or instrumental biases.
keywords:
stellar-mass black hole , accretion disk , X-ray observations , quantum gravity phenomenology , asymptotic safe gravity
journal: Journal of High Energy Astrophysics
[…]
[1]organization=Osservatorio Astronomico di Brera, Istituto Nazionale di Astrofisica (INAF),addressline=Via E. Bianchi, 46, city=Merate, postcode=23807, country=Italy
[…]
[2]organization=Sezione di Milano, Istituto Nazionale di Fisica Nucleare (INFN),addressline=Via G. Celoria, 16, city=Milano, postcode=20133, country=Italy
[…]
[3]organization=Osservatorio Astrofisico di Torino, Istituto Nazionale di Astrofisica (INAF),addressline=Via Osservatorio, 20, city=Pino Torinese, postcode=10025, country=Italy
[…]
[4]organization=Osservatorio Astrofisico di Catania, Istituto Nazionale di Astrofisica (INAF),addressline=Via Santa Sofia, 78, city=Catania, postcode=95123, country=Italy
[…]
[5]organization=Sezione di Catania, Istituto Nazionale di Fisica Nucleare (INFN),addressline=Via Santa Sofia, 64, city=Catania, postcode=95123, country=Italy
{highlights}
We searched for deviations from general relativity on astrophysical scales, particularly close to the innermost stable circular orbit around stellar-mass black holes (calculated according to the Kerr metric).
We selected a number of cases in the available literature, made them homogeneous, and compared with the expectations from the Kerr metric: no deviations within the measurement errors were observed.
We set a lower limit to the $\tilde{\xi}$ parameter of a renormalization-group improved Kerr metric in the low-energy regime (Sánchez, 2024).
@@ 1 Introduction
The description of all physical phenomena in the Universe in one unified theory is the ultimate dream of theoretical physics. Despite of decades of theoretical research (see Ashtekar & Petkov 2014; Armas 2021 for recent reviews), little has been obtained on the observational/experimental side. Any experimental or observational evidence of quantum gravity remains elusive still today (Alves Batista et al., 2023). Today more than ever, it is necessary to find some observational constraint, to try stopping the uncontrolled proliferation of theories. As stressed by Eichhorn & Held (2023), “irrespective of theoretical considerations, any observational avenue to put constraints on deviations from general relativity (GR), should be explored” . This is also our philosophical approach to the problem, searching for any deviation from GR around black holes [^1] In the present work, we always adopted the Kerr metric as reference for GR calculations. Therefore, whether we talk about GR or explicitly write Kerr metric, we mean to refer to Novikov & Thorne (1973)., although we eventually compare observations with expectations of asymptotic safe gravity (ASG), which we consider to be the best option presently at hands.
First developed by Weinberg (1979), ASG has its pillars on quantum field theory and renormalization group (for example, see Niedermaier & Reuter 2006; Bonanno et al. 2020a for reviews). The key point is that quantum effects are described by an unknown parameter $\xi$, which changes the gravitational constant, making it to depends on the energy scale. In the low energy limit (infrared), the running coupling can be written as:
$G(r)=G_{\rm N}\left(1-\frac{\xi}{r^{2}}\right)$ (1)
where $G_{\rm N}=6.67428\times 10^{-8}$ cm3 g-1 s-2 is the classical gravitational constant.
ASG is expected to have observational consequences on astrophysical and cosmological scales (Reuter & Weyer, 2004), and in particular around black holes (Reuter & Tuiran, 2011; Haroon et al., 2018; Eichhorn & Held, 2023; Sánchez, 2024) and neutron stars (Bonanno et al., 2020b). The main effect on black holes is a more compact object, although without curvature singularities (Eichhorn & Held, 2023). A smaller event horizon affects also all the other related quantities such as the photon ring (Eichhorn & Held, 2023), the ergosphere (Haroon et al., 2018) and the radius of the innermost stable circular orbit $r_{\rm isco}$ (Sánchez, 2024). Particularly, Sánchez (2024) suggested that differences up $\sim 30$% in the $r_{\rm isco}$ are expected in the case of maximally rotating prograde black holes already with a small value of $\xi$ (see his Table I, where he adopted $\tilde{\xi}=\xi/r_{\rm g}^{2}$, being $r_{\rm g}$ the gravitational radius of the black hole [^2] Please note that Sánchez (2024) adopted geometrized units with $c=G_{\rm N}=1$. Therefore, in his work $\tilde{\xi}=\xi/\bar{M}^{2}$, where $\bar{M}$ is the mass of the black hole in units of length, i.e. its gravitational radius.).
Some attempts to test these theories have been already made by using X-ray spectroscopy (see Zhang et al. 2018; Zhou et al. 2021; Bambi 2024), but to date no deviations were found. We focus our study on the search for deviations from the value of the $r_{\rm isco}$ in the case of stellar-mass black holes calculated according to Kerr metric (Novikov & Thorne, 1973). Before starting new data analyses and model development, we decided to search in the scientific literature for cases worthy of study.
@@
Stellar-mass black holes are the smallest singularities, as primordial black holes were never observed to date. Therefore, the effects of ASG – if any – should be maximised. Moreover, since the peak of the emission from the accretion disk is inversely proportional to the square root of the mass of the compact object, it falls in the X-ray energy band, where there are less problems of contamination from the nearby environment. In the case of supermassive black holes, the disk emission peaks in the ultraviolet, where there are contaminations from emission lines of the broad-line region and the host galaxy. In the case of advection dominated accretion disks with low accretion rate, which peak at microwave/infrared frequencies (Beckert & Duschl, 2002), it is possible to use the excellent spatial resolution of the Event Horizon Telescope to probe scales $\sim 10r_{\rm g}$ (EHT Collaboration, 2019, 2022). However, $r_{\rm isco}$ can be very close to $r_{\rm g}$ in the case of maximally rotating prograde black hole.
As a first step, we searched in all the available literature for measurements of the inner radius of the accretion disk when the latter was in high/soft state (see Sect. 3). Then, we made all the values homogeneous so we can compare them with each other, and with the theoretical values calculated according to the Kerr metric. Given the measurement errors, we conclude by setting a lower limit to the $\tilde{\xi}$ parameter of a renormalization-group improved Kerr metric in the low-energy regime (Sánchez, 2024).
@@ 2 A short summary of the accretion disk theory and models
According to the standard theory (Shakura & Sunyaev, 1973; Novikov & Thorne, 1973), the accretion disk around black holes is optically thick and geometrically thin, although there are other models, particularly in the case of very high or very low mass accretion rate (Abramowicz & Fragile, 2013). From the observational point of view, the disk can be considered as a series of rings of increasing size going outward, each one emitting black body radiation with increasing temperature $T(r)$ as the radial distance $r$ from the black hole decreases (e.g. see Fig. 2 in Ebisawa et al. 2003). The highest temperature $T_{\rm in}$ corresponds to the inner part of the disk, closer to the event horizon. The most used model in X-ray data analysis is the multicolor accretion disk (Mitsuda et al., 1984; Makishima et al., 1986, 2000), and is implemented as diskbb model in the xspec software package [^3] https://heasarc.gsfc.nasa.gov/xanadu/xspec/. This model is based on a Newtonian approximation, but it can give results consistent with a Kerr black hole under certain inclination angles (small values, so that the Doppler boosting and the gravitational redshift are negligible) and with proper correction coefficients (Shimura & Takahara, 1995; Kubota et al., 1998), as explained in many studies (see, for example, Merloni et al. 2000; Ebisawa et al. 2003). It is worth noting that diskbb does not satisfy the zero-torque condition at the inner boundary. There are two options to cope with this issue: the most adopted and simplest option is to apply a correction factor (Kubota et al., 1998) to the value of the inner radius of the accretion disk as measured by diskbb, but there is also a specific xspec model developed by Zimmerman et al. (2005), namely ezdiskbb, which gives similar results to the previous method since both are based on the same equation (cf Eq. 1 of Kubota et al. 1998 in the case of $\beta=1$, with Eq. 4 of Zimmerman et al. 2005).
There is also a fully relativistic model of multicolor accretion disk (kerrbb in xspec, Li et al. 2005). However, since we are searching in the literature, and diskbb is the most adopted model across many years, we decided to refer to it for our calculations, and to apply the proper correction factors. In addition, when adopted, kerrbb is used to estimate the spin, rather than $r_{\rm isco}$. Therefore, we can make some comparison only in those cases when the parameters $M$ (black hole mass), $i$ (viewing angle), and $d$ (distance), which are generally fixed in the model, are consistent with the values we adopted as reference (see Sect. 4.1).
The diskbb model can be summarised as follows. Although the emitted spectrum is a superposition of several blackbody at different temperatures, the disk luminosity $L_{\rm d}$ [erg s-1] can be calculated via the Stefan-Boltzmann law:
$L_{\rm d}=4\pi r_{\rm in}^{2}\sigma T_{\rm in}^{4}$ (2)
where $\sigma=5.67\times 10^{-5}$ erg cm-2 s-1 K-4 is the Stefan-Boltzmann constant, and $T_{\rm in}$ [K] is the temperature at the inner radius $r_{\rm in}$ [cm]. This equation is valid when the outer radius of the accretion disk $r_{\rm out}$ is much larger than $r_{\rm in}$, so that the outer temperature $T_{\rm out}$ can be considered negligible (Mitsuda et al., 1984; Makishima et al., 1986, 2000).
However, X-ray spectra cannot measure the effective temperature, because of Comptonization and relativistic effects. The observed color temperature $T_{\rm col}$ is greater than the effective one by a factor $\kappa\sim 1.7-2.0$, which depends on the accretion rate (Shimura & Takahara, 1995). This is also linked to the effective location of the inner radius, which should be corrected by a factor $\varsigma=0.412$ (Kubota et al., 1998). Therefore, Eq. (2) is rearranged as:
$L_{\rm d}=4\pi\sigma\left(\frac{r_{\rm in}}{\varsigma}\right)^{2}\left(\frac{T_{\rm col}}{\kappa}\right)^{4}$ (3)
@@
$L_{\rm d}=2\pi\frac{d^{2}}{\cos i}F_{\rm bol}$ (4)
where $F_{\rm bol}$ is the disk bolometric flux [erg cm-2 s-1], $d$ is the luminosity distance [cm], and $i$ is the viewing angle of the disk [deg]. By combining Eqs. (3) and (4), it is then possible to estimate the inner radius of the accretion disk:
$r_{\rm in}=\varsigma\kappa^{2}d\sqrt{\frac{F_{\rm bol}}{2\cos i\sigma T_{\rm col}^{4}}}$ (5)
The normalization $N$ of the diskbb model – that can be directly estimated by spectral fitting – includes some of the above factors, and is equal to:
$N=\left(\frac{r_{\rm col}}{d_{10}}\right)^{2}\cos i$ (6)
where $d_{10}$ is the luminosity distance in units of $10$ kpc, and $r_{\rm col}=r_{\rm in}/\varsigma$ is the color radius. Therefore, the inner radius can also be calculated by rearranging Eq. (6) as:
$r_{\rm in}=\varsigma\kappa^{2}r_{\rm col}=\varsigma\kappa^{2}d_{10}\sqrt{\frac{N}{\cos i}}$ (7)
To conclude, as already noted and with all the caveats (Merloni et al., 2000; Ebisawa et al., 2003), despite its simplicity, the diskbb model performs reasonably well, as it results from the comparison with more detailed relativistic models (e.g. Ebisawa et al. 1991, 2003; Foschini et al. 2006; Lorenzin & Zampieri 2009).
@@ 3 Gathering data
Eqs. (5) and (7) offer two possibilities to estimate the inner radius of the accretion disk. In addition to the quantities measured via the fit of the X-ray spectrum ($N$, $T_{\rm col}$, $F_{\rm bol}$), it is necessary to know the distance $d$, the viewing angle $i$, but also the mass of the black hole $M$ and its spin $a$ to estimate the radius of the innermost stable circular orbit $r_{\rm isco}$ according to GR (Kerr metric, Novikov & Thorne 1973):
$r_{\rm isco}=r_{\rm g}\left[3+Z_{2}-\sqrt{(3-Z_{2})(3+Z_{1}+2Z_{2})}\right]$ (8)
@@
It is worth stressing, however, that the spin is the most critical value to be measured, and the debate is quite hot still today (Fabian et al., 2012; Reynolds & Fabian, 2008; Salvesen & Miller, 2021; Zdziarski et al., 2024a, b; Belczynski et al., 2024; Lasota & Abramowicz, 2024). There are basically two methods: the disk continuum fitting and the iron line modelling. The former is based on the determination of $r_{\rm isco}$, assuming that it is coincident with $r_{\rm in}$, by using diskbb or any other model of accretion disk. Given $r_{\rm g}$ and by using Eq. (8), then it is possible to calculate $a$. In practice, this is just the inverse of the method adopted in this work. Therefore, it is not suitable for our goal.
@@
The values of $r_{\rm g}$, $a$, $d$, and $i$ changed often over time, depending on the technological evolution of the instruments and the observation conditions (e.g. finding the stellar companion is often difficult, given the presence of many molecular clouds and interstellar dust toward the Galactic centre). We have adopted the most recent values, which are not always consistent with those used by the authors we examined in the conversion of their observational data. For example, many authors did not report the normalization $N$ of the diskbb model, but only the calculated $r_{\rm in}$ or even $r_{\rm col}$. Therefore, to extract a value comparable with other observations made at different epochs and with different instruments, it is necessary to know which values of $d$ and $i$ were adopted in the conversion. This is not granted: sometimes the values used in the conversion were not written in the paper, which made it impossible to recalculate the original normalization. Other problems encountered when searching in the literature were: missing measurement errors, missing units of measurement, plain errors and/or typos. In case of errors, it was often possible to correct them by recalculating the described model. For example, when finding a weird normalization $N$, it is possible to recalculate it if $T_{\rm col}$ and $F_{\rm bol}$ were reported. Fluxes were often reported without measurement errors: in this case, a typical $10$% error was adopted.
Yet another issue regards the instruments. In the case of stellar-mass black holes, the color temperature of the multicolor accretion disk peaks in the range $\sim 0.1-1$ keV. However, many detectors have a low-energy threshold of a few keV, which made it difficult to correctly extrapolate to lower energies. For example, one of the most used and prolific satellites in the observation of X-ray binaries was the Rossi X-ray Timing Explorer (RXTE) [^4] https://heasarc.gsfc.nasa.gov/docs/xte/xtegof.html. The Proportional Counter Array (PCA) worked in the $2-60$ keV energy range, but its low-energy threshold degraded up to $3$ keV with the age. Already the $2$ keV threshold makes it difficult to correctly extrapolate a $T_{\rm col}=0.4$ keV, because only a small part of the spectrum can be detected, as shown in Fig. 1.
@@
In addition to the above cited problems, it is necessary to take into account that the accretion disk does not extend always down to the innermost stable circular orbit. Stellar-mass black holes change their activity from high/soft to low/hard states, and vice versa, passing through some intermediate states, depending on the accretion rate (e.g. Esin et al. 1997; Remillard & McClintock 2006; Belloni 2010; Motta et al. 2021). High/soft states occur at high accretion rates. The X-ray spectrum is dominated by the thermal emission of a geometrically-thin optically-thick accretion disk, which can extend down to $r_{\rm isco}$, and it accounts for more than $50$% of the bolometric flux (Remillard & McClintock, 2006). At low accretion rates (low/hard states), the thin disk is weak and cool, and a mildly relativistic radio jet is present. The X-ray spectrum is dominated by the optically-thin emission from the corona, modelled with a hard power law or a more detailed Comptonization model. The intermediate state is an obvious mixture of the two above cited states. There are also many other variations, distinguished by small details [^5] See the plot: https://www.issibern.ch/teams/proaccretion/Documents.html. However, for our purposes, we focus our attention on high/soft states, when the accretion disk extends downward the innermost stable circular orbit (i.e. $r_{\rm in}\sim r_{\rm isco}$). This is also a requirement to have a reliable measurement of $r_{\rm in}$, because the multicolor disk model tends to underestimate the inner radius when the corona significantly contributes to the bolometric flux (Merloni et al., 2000), as it happens in the intermediate states, for example (see Sect. 4.3).
When many measurements are available in the same article (X-ray spectra fit with different models, multiple observations because of a monitoring campaign), we selected the most promising cases, i.e. the best fit and the smallest values of $r_{\rm in}$.